Metal Works Inc makes a wide variety of heavy duty metal ornaments that are highly sort after

in the home gardens. The company has done very well in recent years, and has just committed to

paying out a $2 dividend on each of their outstanding 50,000 shares. Building on recent success

the company is considering diversifying into the production of lighter and more delicate

ornaments for inside the home, however the company recognises that to do so will require the

purchase of new machinery. Prior to making such a big commitment to pursue this new market

they spent $90,000 on consultants whose considered opinion is that there is likely to be strong

market demand for the company’s new product. The consultants believe that the additional

machinery would cost the firm $750,000 and likely require an increase of $65,000 in raw materials

held on site. These additional materials would be required during the operating life of the

machinery. Further, they estimated pre-tax cash-flows of $220,000 per year and a likely salvage

value of $280,000 when the asset is sold in 8 years. The company has a cost of capital of 12% and

a tax rate of 30%. For any depreciation use a CCA rate 30% and assume the asset class will remain

open and is well-funded (thus salvage value< UCC of the class).

a) What is the initial cash outlay at t=0 ?

b) After one year what value will the asset have depreciated to under the CCA rules?

c) What is the ending cash flow at time t=8?

d) What is the NPV of the project? Should the company accept or reject this project?
Answer #1
NOTE:
$90,000 spent on consultants is sunk cost and not relevant for this decision
a) Initial Cash Flow at t=0
Cost of additional machinery $750,000
Increase in raw materials $65,000
Initial Cash Flow at t=0 ($815,000)
b) Annual depreciation
Year                                1 2                              3 4                               5 6                              7 8
A Book Value at beginning of Year $750,000 $525,000 $367,500 $257,250 $180,075 $126,053 $88,237 $61,766
B=A*30% Annual depreciation $225,000 $157,500 $110,250 $77,175 $54,023 $37,816 $26,471 $18,530
C=A-B Book Value at END of the year $525,000 $367,500 $257,250 $180,075 $126,053 $88,237 $61,766 $43,236
D=B*Tax Rate(30%) Depreciation Tax Shield $67,500 $47,250 $33,075 $23,153 $16,207 $11,345 $7,941 $5,559
After one year Value of the asset $525,000
Book Value at end of year8 $43,236
Before tax Salvage Value $280,000
Gain on Salvage $236,764 (28000-43236)
Tax on Gain =30%*236764= $71,029
After Tax cash flow from Salvage $208,971 (280000-71029)
c Ending Cash Flow at time t=8
After Tax cash flow from Salvage $208,971
Release of initial inventory increase $65,000
Ending Cash Flow at time t=8 $273,971 (208971+65000)
Present Value(PV) of Cash Flow:
(Cash Flow)/((1+i)^N)
i=discount rate=Cost of Capital=12%=0.12
N=Year   of Cash Flow
N Year 0 1 2 3 4 5 6 7 8
A Initial Cash Flow -$815,000
B Pretax Cash Flows $220,000 $220,000 $220,000 $220,000 $220,000 $220,000 $220,000 $220,000
C=B*(1-0.3) After Tax Cash Flow $154,000 $154,000 $154,000 $154,000 $154,000 $154,000 $154,000 $154,000
D Depreciation Tax shield $67,500 $47,250 $33,075 $23,153 $16,207 $11,345 $7,941 $5,559
E=C+D Annual Operating Cash Flow $221,500 $201,250 $187,075 $177,153 $170,207 $165,345 $161,941 $159,559
F Ending Cash Flow at time t=8 $273,971
CF=A+E+F Net Cash Flow -$815,000 $221,500 $201,250 $187,075 $177,153 $170,207 $165,345 $161,941 $433,530 SUM
PV=CF/(1.12^N) Present Valure -$815,000 $197,768 $160,435 $133,156 $112,584 $96,580 $83,769 $73,254 $175,095 $217,641
NPV=Sum of PV Net Present Value(NPV) $217,641
Company Should Accept the project
NPV is Positive